OpenStudy (kainui):

The remainder theorem states that for any polynomial f(x) divided by (x-a) the remainder is always f(a). How do I prove this?

OpenStudy (anonymous):

It is advisable to use: \[\color{green}{\textsf{Dividend = Divisor} \times \textsf{Quotient + Remainder}}\]

OpenStudy (anonymous):

Here, f(x) is your Dividend, Suppose you divide this polynomial f(x) by (x-a), and suppose you get quotient as : q(x), some other polynomial with Remainder R, then: \[f(x) = (x-a) \times q(x) + R\]

OpenStudy (kainui):

so now if we plug in x=a we get f(a)=R ok cool. But now I notice something really interesting.

OpenStudy (anonymous):

What is that??

OpenStudy (kainui):

\[\LARGE f(x)=q(x)(x-a)+f(a)\] if we rearrange this we can get: \[\LARGE \frac{f(x)-f(a)}{x-a}=q(x)\] which looks like the mean value theorem or something. So what's the connection going on here?

OpenStudy (anonymous):

Remainder is f(a) only for x = a..

ganeshie8 (ganeshie8):

all it says is (x-a) divides f(x) evenly when we subtract f(a) from f(x)

OpenStudy (anonymous):

limit definition :)